# Comparison of lubrication theory and Stokes flow models in step bearings with flow separation

### v1

@@ Abstract
The Reynolds equation from lubrication theory and the Stokes equations for low Reynolds number flows are distinct models for an incompressible fluid with negligible inertia. Here we investigate the sensitivity of the Reynolds equation to large gradients in the surface geometry. We present an analytic solution to the Reynolds equation in a piecewise-linear domain alongside a more general finite difference solution. For the Stokes equations, we use a finite difference solution for the biharmonic stream-velocity formulation. We compare the fluid velocity, pressure, and resistance for various step bearing geometries in the lubrication and Stokes limits. We find that the solutions to the Reynolds equation do not capture flow separation resulting from large cross-film pressure gradients. Flow separation and corner flow recirculation in step bearings are explored further; we consider the effect of smoothing large gradients in the surface geometry in order to recover limits under which the lubrication and Stokes approximations converge.
@@ 2.1 Boundary conditions
We consider solutions to Reynolds equation in the domain $[x_{0},x_{L}]\times[y_{h}(x),y_{H}]$ subject to the no-slip fluid boundary condition. The film height is $h(x)=y_{H}-y_{h}(x)>0$. We assume, without loss of generality, that the relative velocity $\mathcal{U}$ is imposed at the upper surface $y=y_{H}$ and the lower surface $y=y_{h}(x)$ is stationary,
@@
[…]
@@ 2.2.1 Finite difference solution
Define a uniform discretisation of $[x_{0},x_{L}]$,
@@ 2.2.2 Analytic solution
(20) $\lim_{x\to x_{k}^{-}}p(x)$ =$=\begin{cases}\frac{-1}{G_{k-1}}\Big{(}\frac{1}{2}\mathcal{C}_{Q}h_{k^{-}}^{-2}+6\eta\mathcal{U}h_{k^{-}}^{-1}\Big{)}+\mathcal{C}_{P_{k-1}}&G_{k-1}\neq 0\\ (x_{k}-x_{k-1})\Big{(}\mathcal{C}_{Q}h_{k^{-}}^{-3}+6\eta\mathcal{U}h_{k^{-}}^{-2}\Big{)}+\mathcal{C}_{P_{k-1}}&G_{k-1}=0\end{cases},$
(21) $\lim_{x\to x_{k}^{+}}p(x)$ =$=\begin{cases}\frac{-1}{G_{k}}\Big{(}\frac{1}{2}\mathcal{C}_{Q}h_{k^{+}}^{-2}+6\eta\mathcal{U}h_{k^{+}}^{-1}\Big{)}+\mathcal{C}_{P_{k}}&G_{k}\neq 0\\ \mathcal{C}_{P_{k}}&G_{k}=0\end{cases}.$
@@ 3 The biharmonic Navier-Stokes equation
The stream-velocity formulation of the incompressible Navier-Stokes equations is a method of solution for low to moderate Reynolds number flows [19], [9], [18], [11], [15], [4].
@@
are expressed in stream-velocity form,
@@
The Reynolds number $Re=\mathcal{U}L_{y}/\nu$ controls the inertial terms. When $Re=0$, the stream-velocity formulation Eq. 27 reduces to the biharmonic Stokes equation $\nabla^{4}\psi=0$.
@@ 3.1 Boundary conditions
We consider solutions in the domain $[x_{0},x_{L}]\times[y_{h}(x),y_{H}]$, where $h(x)=y_{H}-y_{h}(x)>0$. The no-slip boundary condition is imposed at the upper and lower surfaces. The surface boundary velocities are prescribed,
@@ 3.2 Finite difference iterative solution
Define the uniform discretisation of the domain $[x_{0},x_{L}]\times[\min y_{h}(x),y_{H}]$,
@@
The stream-velocity definition Eq. 28 is discretised as,
@@
The equations Eqs. 43, 44 and 45 are evaluated iteratively. The biharmonic equation Eq. 43 is solved for $\psi$, then the stream-velocity relationship Eqs. 44 and 45 updates $u$ and $v$. The initial values for $\psi$, $u$ and $v$ satisfy the boundary conditions and are otherwise zero. The iterative process is terminated when the maximum absolute error in $\psi$ between subsequent iterations falls below $10^{-8}$. This cut off point for convergence is taken in the literature [9], [4]. If the bulk flow has $\psi$ of unit magnitude, regions of $|\psi|\leq 10^{-8}$ are found very close to the corner [13], and are not distinguishable by grid resolutions at our disposal. As a consequence, although in principle there is an infinite sequence of Moffatt eddies to be found in corner geometries, the convergence cut off point and the grid resolution set a limit on the number of eddies that are actually resolved in practice.
@@ 3.3 Pressure from velocity
The pressure is then determined by a path integral from the outlet. We expect path independence for a stream-velocity solution that has sufficiently converged. For convenience, the contour is taken first from outlet to inlet along the upper surface,
@@ 4 Comparing the Reynolds and Stokes solutions
[…]
@@ 4.1 The backward facing step
Flow separation in the BFS is assessed for various expansion ratios $\mathcal{H}$. The corner separated region is characterized by the separation point $(L_{\text{in}},y_{r})$ and the reattachment point $(L_{\text{in}}+x_{r},0)$. As shown in Fig. 3, the separation and reattachment lengths $y_{r}$ and $x_{r}$ increase steadily with increasing expansion ratio. Close-up views of the corner recirculation in the BFS are shown in Fig. 4. Secondary recirculation zones are captured on zooming closer in to the corner. The size and intensity of each sequential eddy decreases rapidly on approach to the corner.
@@ 4.3 A smoothed step bearing
The Reynolds and Stokes solutions for the smoothed step bearing are evaluated for varying $20\leq\delta\ll L/$ 2, and a range of expansion ratios $1<\mathcal{H}\ll L$. Results are presented for various $H_{\text{out}}$ and constant $H_{\text{in}}=1$, $L_{\text{in}}=1$, $L=4$. We analyze the error in fluid resistance $\mathcal{R}$ for the Reynolds and Stokes solutions relative to the slope reciprocal $\delta$ and the expansion ratio $\mathcal{H}$.
@@ 5 Conclusions
Various examples of step bearings were observed with the Reynolds and Stokes equations. We considered two methods of solution to the Reynolds equation, Section 2. For textured bearings which are piecewise-linear by nature, an analytic solution is available in addition to the more general numerical finite difference method. The Stokes equations were solved using an iterative finite difference method for the biharmonic stream-velocity formulation, Section 3. The discrepancies observed between solutions to the Reynolds and Stokes equations characterize the extent to which the lubrication assumptions are applicable for each geometry.

### v2

@@ Abstract
The Reynolds equation from lubrication theory and the Stokes equations for low Reynolds number flows are distinct models of an incompressible fluid with negligible inertia. Here we investigate the sensitivity of the Reynolds equation to large gradients in the surface geometry. We present an analytic solution to the Reynolds equation in a piecewise-linear domain alongside a more general finite difference solution. For the Stokes equations, we use a finite difference solution for the biharmonic velocity-stream formulation. We compare the fluid velocity, pressure, and resistance for various step bearing geometries in the lubrication and Stokes limits. We find that the solutions to the Reynolds equation do not capture flow separation resulting from large cross-film pressure gradients. Flow separation and corner flow recirculation in step bearings are explored further; we consider the effect of smoothing large gradients in the surface geometry in order to recover limits under which the lubrication and Stokes approximations converge.
@@ 2.1 Boundary conditions
We consider solutions to Reynolds equation in the domain $x_{0},x_{L}]\times[y_{h}(x),y_{H}]$ subject to the no-slip fluid boundary condition. The film height is $h(x)=y_{H}-y_{h}(x)>0$. We assume, without loss of generality, that the relative velocity $\mathcal{U}$ is imposed at the upper surface $y=y_{H}$ and the lower surface $y=y_{h}(x)$ is stationary,
@@
(10) $\mathcal{Q}=\int_{y_{h}(x)}^{y_{H}}u(x,y)\,dy=\frac{-1}{12\eta}\Bigg(\Big[h(x)\Big]^{3}\frac{dp}{dx}-6\eta\mathcal{U}h(x)\Bigg).$
@@ 2.2.1 Finite difference solution
Define a uniform discretisation of $x_{0},x_{L}]$,
@@ 2.2.2 Analytic solution
(20) $\lim_{x\to x_{k}^{-}}p(x)$ $=\begin{cases}\frac{-1}{G_{k-1}}\Big(\frac{1}{2}\mathcal{C}_{Q}h_{k^{-}}^{-2}+6\eta\mathcal{U}h_{k^{-}}^{-1}\Big)+\mathcal{C}_{P_{k-1}}&G_{k-1}\neq 0\\ (x_{k}-x_{k-1})\Big(\mathcal{C}_{Q}h_{k^{-}}^{-3}+6\eta\mathcal{U}h_{k^{-}}^{-2}\Big)+\mathcal{C}_{P_{k-1}}&G_{k-1}=0\end{cases},$
(21) $\lim_{x\to x_{k}^{+}}p(x)$ $=\begin{cases}\frac{-1}{G_{k}}\Big(\frac{1}{2}\mathcal{C}_{Q}h_{k^{+}}^{-2}+6\eta\mathcal{U}h_{k^{+}}^{-1}\Big)+\mathcal{C}_{P_{k}}&G_{k}\neq 0\\ \mathcal{C}_{P_{k}}&G_{k}=0\end{cases}.$
@@ 3 The biharmonic Navier-Stokes equation
The velocity-stream formulation of the incompressible Navier-Stokes equations is a method of solution for low to moderate Reynolds number flows [19], [9], [18], [11], [15], [4].
@@
are expressed in velocity-stream form,
@@
The Reynolds number $Re=\mathcal{U}L_{y}/\nu$ controls the inertial terms. When $Re=0$, the velocity-stream formulation Eq. 27 reduces to the biharmonic Stokes equation $\nabla^{4}\psi=0$.
@@ 3.1 Boundary conditions
We consider solutions in the domain $x_{0},x_{L}]\times[y_{h}(x),y_{H}]$, where $h(x)=y_{H}-y_{h}(x)>0$. The no-slip boundary condition is imposed at the upper and lower surfaces. The surface boundary velocities are prescribed,
@@ 3.2 Finite difference iterative solution
Define the uniform discretisation of the domain $x_{0},x_{L}]\times[\min y_{h}(x),y_{H}]$,
@@
The velocity-stream definition Eq. 28 is discretised as,
@@
The equations Eqs. 43, 44 and 45 are evaluated iteratively. The biharmonic equation Eq. 43 is solved for $\psi$, then the velocity-stream relationship Eqs. 44 and 45 updates $u$ and $v$. The initial values for $\psi$, $u$ and $v$ satisfy the boundary conditions and are otherwise zero. The iterative process is terminated when the maximum absolute error in $\psi$ between subsequent iterations falls below $10^{-8}$. This cut off point for convergence is taken in the literature [9], [4]. If the bulk flow has $\psi$ of unit magnitude, regions of $|\psi|\leq 10^{-8}$ are found very close to the corner [13], and are not distinguishable by grid resolutions at our disposal. As a consequence, although in principle there is an infinite sequence of Moffatt eddies to be found in corner geometries, the convergence cut off point and the grid resolution set a limit on the number of eddies that are actually resolved in practice.
@@ 3.3 Pressure from velocity
The pressure is then determined by a path integral from the outlet. We expect path independence for a velocity-stream solution that has sufficiently converged. For convenience, the contour is taken first from outlet to inlet along the upper surface,
@@ 4 Comparing the Reynolds and Stokes solutions
(53) $\mathcal{R}=\frac{\Delta\mathcal{P}}{-\mathcal{Q}}=\frac{p(x_{L},y)-p(x_{0},y)}{\int_{y_{h}}^{y_{H}}u(x,y)\,dy}.$
@@ 4.1 The backward facing step
Flow separation in the BFS is assessed for various expansion ratios $\mathcal{H}$. The corner separated region is characterized by the separation point $(L_{\text{in}},y_{r})$ and the reattachment point $(L_{\text{in}}+x_{r},0)$. As shown in Fig. 3, the separation and reattachment lengths $y_{r}$ and $x_{r}$ increase steadily with increasing expansion ratio. The reattachment length $x_{r}/h=0.3562$ for $H/h=2$ is consistent with previously reported results in the literature. Biswas et al. [3] analyze the BFS at a range of Reynolds numbers for fixed expansion ratio. They observe that the reattachment length for expansion ratio $H/h=1.9423$ approaches a constant size $x_{r}/h\to 0.35$ in zero Reynolds number limit. Close-up views of the corner recirculation in the BFS are shown in Fig. 4. Secondary recirculation zones are captured on zooming closer in to the corner. The size and intensity of each sequential eddy decreases rapidly on approach to the corner.
@@ 4.3 A smoothed step bearing
The Reynolds and Stokes solutions for the smoothed step bearing are evaluated for varying $0\leq\delta\ll L/2$, and a range of expansion ratios $1<\mathcal{H}\ll L$. Results are presented for various $H_{\text{out}}$ and constant $H_{\text{in}}=1$, $L_{\text{in}}=1$, $L=4$. We analyze the error in fluid resistance $\mathcal{R}$ for the Reynolds and Stokes solutions relative to the slope reciprocal $\delta$ and the expansion ratio $\mathcal{H}$.
@@ 5 Conclusions
Various examples of step bearings were observed with the Reynolds and Stokes equations. We considered two methods of solution to the Reynolds equation, Section 2. For textured bearings which are piecewise-linear by nature, an analytic solution is available in addition to the more general numerical finite difference method. The Stokes equations were solved using an iterative finite difference method for the biharmonic velocity-stream formulation, Section 3. The discrepancies observed between solutions to the Reynolds and Stokes equations characterize the extent to which the lubrication assumptions are applicable for each geometry.